An electron is traveling horizontally at a speed of 3.75x10^6 m/s and enters in a uniform electric field. a) What is the magnitude and direction of the electric à ƒ ƒ ¯ ¬ eld that uniformly stops the electron in a distance of 28 mm? b) How much time does it take to the electron to stop? Hint: Use the kinematic equations and Newtonà ƒ ƒ ¢ € ™s laws.
1
Expert's answer
2021-10-13T15:42:26-0400

a) Negative direction.


"qEd=0.5mv^2\\(1.75\cdot10^{11})(0.028)E=0.5(3.75\cdot10^{6})^2\\E=1435\frac{N}{C}"


b)

"\\t=2\frac{0.028}{3.75\cdot10^{6}}=1.49\cdot10^{-8}\ s"